OpenStudy (mathhater):

-5 minus 3(x minus 2) = -2

OpenStudy (anonymous):

nope, I guess google says the answer is 1, not sure what I did wrong

OpenStudy (anonymous):

-5 - 3(x-2)=-2 -5 - (3x+6=-2 -5 - 3x+ 6=-2 -5-3x= -8 -3x= -3 x= -1